OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point-slope form. Passing through (-5, -7) and (-8, -6) I honestly cant do these, they confuse me

OpenStudy (anonymous):

I have multiple choices , I just don't know:(

OpenStudy (anonymous):

A. y - 7 = -1/3 (x - 5) or y - 6 = -1/3 (x - 8) B. y + 7 = -1/3 (x + 8) or y + 6 = -1/3 (x + 5) C. y + 7 = -1/3 (x + 5) or y + 6 = -1/3 (x + 8) D. y + 7 = -1/3 x - 5 or y + 6 = -1/3 x + 7

OpenStudy (anonymous):

@fateal

OpenStudy (fateal):

http://www.mathpapa.com/algebra-calculator.html i used this link

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

point slope equation (y-y1) = m(x-x1) (x,y) any point (x1,y1) point on the line m= slope

OpenStudy (anonymous):

I got c as my answer

OpenStudy (triciaal):

slope m = y2 -y1/x2-x1 the 2 given point (x1, y1) and (x2, y2)

OpenStudy (anonymous):

(-5,-7) (-8,-6)

OpenStudy (anonymous):

-6- -7/-8 - -5?

OpenStudy (triciaal):

(-5, -7) and (-8, -6) m = -6 --7/-8--5 = -6 +7/(-8 + 5) = 1/-3 = -1/3 y --6=-1/3(x --8) y + 6 = -1/3x + 8/3 y = -1/3x +8/3-18/3 =-1/3x -10/3 point-slope form yes c

OpenStudy (triciaal):

sorry here are the steps 2 points on the line find the slope slope = change in y / CORRESPONDING change in x (y2-y1)/(x2-x1) using any one of the given points find the equation for any point (x, y) equation (y-y1)= m(x-x1) where (x1,y1) is a point on the line same line if you use (x2, y2) any questions?